Groleau Corporation has an activity-based costing system with three activity cost pools--Processing, Setting Up, and Other. The company's overhead costs, which consist of factory utilities and indirect labor, are allocated to the cost pools in proportion to the activity cost pools' consumption of resources. Costs in the Processing cost pool are assigned to products based on machine-hours (MHs) and costs in the Setting Up cost pool are assigned to products based on the number of batches. Costs in the Other cost pool are not assigned to products. Data concerning the two products and the company's costs and activity-based costing system appear below:

Factory utilities (total) $ 29,200
Indirect labor (total) $ 9,500

Distribution of Resource Consumption Across Activity Cost Pools

Processing Setting Up Other
Factory utilities 0.30 0.30 0.40
Indirect labor 0.30 0.50 0.20
MHs Batches
Product S8 3,200 1,100
Product F1 7,200 500
Total 10,400 1,600
Product S8 Product F1
Sales (total) $ 66,200 $ 94,100
Direct materials (total) $ 19,800 $ 36,600
Direct labor (total) $ 26,900 $ 43,800

Required:

a. Assign overhead costs to activity cost pools using activity-based costing.

b. Calculate activity rates for each activity cost pool using activity-based costing.

c. Determine the amount of overhead cost that would be assigned to each product using activity-based costing.

d. Determine the product margins for each product using activity-based costing.
Answer #1
a
Processing Setting Up Other Total
Factory utilities 8760 8760 11680 29200
Indirect labor 2850 4750 1900 9500
Total 11610 13510 13580 38700
b
Processing 1.12 per MH =11610/10400
Setting Up 8.44 per Batch =13510/1600
c
Amount of
overhead costs
Product S8 12868 =(3200*1.12)+(1100*8.44)
Product F1 12284 =(7200*1.12)+(500*8.44)
d
Product S8 Product F1
Sales 66200 94100
Less: Costs
Direct materials 19800 36600
Direct labor 26900 43800
Overhead 12868 12284
Total costs 59568 92684
Product margin 6632 1416